INAF Science with DX AO

04:53 Done with LUCI2 SL integration
  • power on DX WUnit
  • wait to power on because stage homing moves into N3.75 FOV

04:58 Powered on and fastlink done
  • ready for Preset
  • Steve is pointing and collimating in parallel

04:59 PresetAO
  • 12.28 mag
  • apply flat bin1_400modes_26Nov_tilt_removed2.sav
  • Should make this the default
  • 11.7 measured
  • bin 1, 800 Hz, 153 modes (default was 625 Hz)
  • CmdDistrub (625 Hz) 20170127_062504
  • 1.62, .2, 0.38 (change to 1.47, 0.35, 0.33)
  • TN 051825

05:30 Loop closed.
  • forced it to use bin1, 625 Hz (check flux changed it to 800 Hz: Troubles at Act 455 with high forces and rolled edge)
  • forced by no doing Check Flux step

05:31 When observer did continue of LUCI2 script, LUCI sent new preset, even though the previous preset was successful
  • Skipped Check Flux step again
  • set to 1.5, 0.35, 0.3
  • Gopt = 0.4 -> 0.49
  • Luci collect images

05:40 Reference star is elongated
  • TN 054255
  • Find N30 mask is WAY off. Could introduce astigmatism
  • Tried to use AFC tab to align. No motion
  • Called Dave T. He aligned it by hand
  • AO Ref now in center of image

06:03 Resume LUCI Script
  • Still a bit of elongation in luci2.20171127.0170.fits
  • LBTplot show Tip corrections of 24, 25 and Tilt of 18, 35 (usually ~10-15)
  • TN 060903 (50 mas), 061915 (45 mas)
  • See 60 Hz vibration at ~50 mas
  • Jitter Plot with 60 Hz @ 50 mas. See elongation on LIUCI2 image:
    jitter_60Hz_50mas.jpg

06:34 Steve close vent doors
  • TN 063450 (40 mas)
  • not any change (or small) now see 40 was compared to 45 mas 20 minutes ago and 50 mas an hour ago
  • TN 045357 (38 mas)
    • modal plot not looking good
    • modal plot (not too good):
      modalplot_1.jpg

07:12 Optimize Gain
  • 1.7, 0.58, 0.2
    • 071316
  • gopt = 0.35 0.7, 0.18, 0.08
    • 071405
  • 0.35, 0.18, 0.08
    • 071722
  • LUCI 215 image looks a bit better. Start to see a bit of a diffraction ring
  • directly into the wind now

07:46 Loop Opened
  • may have been seeing got much better quickly and gains were too high
  • click re-close loop, optimize gain, gopt, NCPA on
  • 1.3, 0.38, 0.38
  • gopt=0.43 (0.57, 0.17, 0.17)
  • Start LUCI script again

08:04 Preset to PSF star
  • 12.7 mag
  • see 13.7 mag. Change to bin 3. Need bin 2

08:06 Preset to 9.9 mag
  • see 10.6
  • use bin 2, 990 Hz
  • force bin2, 625 Hz
  • 2.25, 0.62, 0.57 (2.0)
    • TN 081435
  • gopt=0.34 0.74, 0.23, 0.21
    • TN 081619
    • Reasonable T/T gain

08:26 Preset to next object
  • 13 mag
  • measure 13.3
  • bin 3, 400 Hz
  • 2.0, 1.0, 1.0
  • no gopt
  • TN 084637

-- %USERSIG{DougMiller - 2017-11-27}%
I Attachment Action Size Date Who Comment
jitter_60Hz_50mas.jpgjpg jitter_60Hz_50mas.jpg manage 62 K 27 Nov 2017 - 06:42 DougMiller Jitter Plot with 60 Hz @ 50 mas. See elongation on LIUCI2 image
modalplot_1.jpgjpg modalplot_1.jpg manage 42 K 27 Nov 2017 - 07:09 DougMiller modal plot (not too good)
Topic revision: r3 - 27 Nov 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback